OpenStudy (anonymous):

What is the slope of the line passing through the point (3, -4) and perpendicular to the line y = 3x + 5? Write your answer as a fraction.

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

if slope of a line is m as in the form y = mx + b then the slope of a line perpendicular to it is -1/m here im = 3 so the answer is -1/3

OpenStudy (anonymous):

y = 3x + 5 Perpendicular slope : m = -1/3 ( there's no need to check the point)